# arXiv:2603.01543v1[math.DG] 02 Mar 2026

### v1

@@ Abstract.
In this paper, we introduce new mass-type invariants for time-symmetric initial data in space-times obeying the Dominant Energy Condition. When the cosmological constant is positive, these invariants, unlike the total Hawking mass, turn out to be genuinely effective in providing new characterizations of de Sitter solution. From a theoretical standpoint, this opens a new perspective on how one might refine the rigidity statement originally proposed by Min-Oo in his well known conjecture, later refuted by the counterexamples of Brendle, Marques, and Neves.
@@
Keywords: Monotonicity formulas, $p$-Green’s function, Hawking mass, de Sitter solution, Min-Oo conjecture.
@@
The geometric intuition behind this statement is quite natural: if one postulates the validity of a positive mass theorem and assumes a strong asymptotic decay of the metric towards the model one – so that any reasonable asymptotic geometric invariant, including the yet undefined notion of mass, is forced to vanish – then it would be natural to conclude that the given manifold is isometric to the model space. In this sense, Theorem 1.2 should be regarded as a pre-Positive Mass Theorem, as it represents a preliminary condition for the validity of a full Positive Mass statement. Following a similar rationale, Min-Oo proposed in [54] his famous conjecture for initial data with a positive cosmological constant, aiming to characterize the de Sitter solution as the only one with zero mass. The precise statement of the conjecture is as follows:
@@
The general conjecture, however, does not hold under the mere natural assumptions under which it was originally formulated: Brendle, Marques and Neves [22] constructed smooth metrics on the upper hemisphere that coincide with the standard round metric in a tubular neighborhood of the equatorial boundary, having scalar curvature ${\mathrm{R}}_{g}\geq 2\Lambda>0$ everywhere and strictly grater than $2\Lambda 0$ on an open domain (supported far away form the round tubular neighborhood of the equator). The particular features of these counterexamples not only disprove the conjecture but also invalidate the heuristic behind it: no pre-positive mass theorem in the style of Theorem 1.2 can be obtained by imposing any order of contact with the model solution – here, the hemisphere or space-like de Sitter metric – in a neighborhood of the cosmological horizon. In physical terms, one may assert that whatever the mass associated with these initial data may be, it cannot be measured by an observer on the cosmological horizon, even under the idealized assumption that this observer possesses complete knowledge of the surroundings.
@@
We now ask whether this strategy can be implemented on a compact initial data set $(M,g)$ with connected minimal boundary and scalar curvature ${\mathrm{R}}\geq 2\Lambda>0$, in line with condition (1.2). Assume the above monotonicity formula extend rigorously to a weak setting that allows flow singularities, in the spirit of Huisken-Ilmanen [40]. Further assume there exists a weak IMCF starting from a point of the initial data set – more precisely, an ancient solution whose level sets collapse to a chosen pole as $t\to$-$\infty$ – and that this flow expires by landing on the cosmological horizon in finite time $T^{*}$. In this thought experiment, the Hawking mass at time $T^{*}$, can be viewed as the total mass of the manifold. Monotonicity, together with $\mathfrak{m}_{\rm Haw}(\Sigma_{t})\to 0$ as $t\to$-$\infty$, then strongly suggests a Positive Mass–type statement for this total mass. This appealing line of reasoning is nevertheless ruled out by the counterexamples of Brendle, Marques, and Neves. Once the flow enters a region where ${\mathrm{R}}>2\Lambda$, the last term in (1.4) forces the derivative of the Hawking mass to be strictly positive. Thus the Hawking mass, which vanishes at the pole and is nondecreasing along the flow, becomes strictly positive and remains so up to time $T^{*}$, when the flow expires on the cosmological horizon. In particular,
@@
However, in these counterexample manifolds a whole neighborhood of the boundary is isometric to a tubular neighborhood of the equator in the standard hemisphere of radius $R_{\Lambda}=\sqrt{3/\Lambda}$ (equivalently, with constant sectional curvature $\Lambda/3$). A direct computation shows that the Hawking mass of this standard equator vanishes, since it is a minimal surface of area $4\pi R_{\Lambda}^{2}=12\pi/\Lambda$. Consequently, any attempt to obtain a Positive Mass-type result via such an inverse mean curvature flow faces fundamental obstructions.
Upon closer examination, it becomes clear that the underlying cause of this speculative contradiction lies in the initial assumption that a flow of the specified type actually exists. Recent progress, including the important work of Xu [73], shows that the Dirichlet problem is not the most appropriate setting in which to study the weak IMCF on a manifold with boundary. Instead, the natural condition appears to be that of the so‑called exterior obstacle condition, in which the level sets of the arrival time function of the IMCF, namely, the weak solution of
@@
where $\nu$ denotes the outward unit normal to $\partial M$ (we refeer the reader to [73, Summary of Definitions/Theorem 1.5] for the precise definition). The Hawking mass, however, fails to be monotone along this type of flow in general. A concrete counterexample comes from the interior of a Clifford torus in the round sphere of radius $R_{\Lambda}$. This manifold has scalar curvature constantly equal to $2\Lambda>0$, so it satisfies (1.2), and its boundary is a minimal Clifford torus $\mathbb{T}_{\Lambda}$ with area $6\pi^{2}/\Lambda$. The Hawking mass of this boundary is therefore
@@
Now start the IMCF with obstacle from a point in this manifold. As noted above, we have $\mathfrak{m}_{\rm Haw}(\Sigma_{t})\to 0$ as $t\to$-$\infty$. However, after the first contact time with the boundary, the Hawking masses of the evolving surfaces $\Sigma_{t}$ are driven closer and closer to the Hawking mass of the boundary. In fact, for times $t$ sufficiently close to the last contact time $T^{*}$ – which is finite, as follows from [73, Theorem 1.7]) – $\mathfrak{m}_{\rm Haw}(\Sigma_{t})$ is strictly negative. Thus, in this example $t\mapsto\mathfrak{m}_{\rm Haw}(\Sigma_{t})$ starts near 0 as $t\to$-$\infty$ and becomes negative as $t\to T^{*}$, so it cannot be nondecreasing for all times.
1.3. A potential-theoretic approach: statement of the main results.
@@
$K_{p}\,=\,\frac{\Gamma(\frac{1}{2})\,\Gamma(c_{p})}{\Gamma(a_{p}+\frac{3}{2})\Gamma(b_{p}+\frac{3}{2})}\,$,
@@
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with compact smooth boundary $\partial M$, whose scalar curvature $\mathrm{R}$} satisfies
@@
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth minimal boundary $\partial M$, whose scalar curvature $\mathrm{R}$} satisfies
@@
in the sense described by Xu in [73, Summary of Definitions/Theorem 1.5], $\nu(y_{0})$ being the outer unit normal at $y_{0}\in\partial M$. The normalization constant in front of the Dirac delta is reflected in the fact that the integral identity (3.30), valid for every $t\in\mathbb{R}$,
@@
induces the following normalization for the area growth law of the IMCF
@@
Note that $T_{\Lambda}$ coincides both with the maximal existence time and with the first contact time of the model IMCF, namely the IMCF starting from the north pole of the standard sphere of radius $R_{\Lambda}$ and having the equator as an outer obstacle. The behavior of the structural coefficients, as $p\to 1^{+}$, makes it possible to identify a formal limit of the monotone quantities toward the Hawking mass. In fact, assuming that the pp‑IMCF $w_{x}^{(p)}$ converge to an IMCF with obstacle $w_{x}^{(1)}$ in a way that is strong enough to justify the convergence of the areas of the level sets and of suitable integral norms of the gradients, we would obtain the following limit for small enough $t<T_{\Lambda}$:
@@
$T_{*}(x)\,=\,{\rm ess}\!\!\!\inf_{\partial M\quad}\!w_{x}^{(1)}\,=\,\inf\left\{t\in\mathbb{R}\,:\,|\partial^{*}\Omega_{x,\tau}^{(1)}\,\cap\partial M|\neq 0\right\}$ (1.8)
of the weak IMCF with the obstacle $\partial M$ emanating from $x$. In this range, in fact, the standard area growth $|\Sigma_{x,t}^{(1)}|=4\pi e^{t}$ is satisfied (see [73, Remark 3.5-(iii)]) and it is very plausible that one can adapt the theory of improved convergence, recently developed by Benatti, Pluda, and Pozzetta in [14, Theorem 1.2].
On one side, the formal convergence to the Hawking mass for fixed times clarifies the geometric nature of our monotone quantities. On the other, it is already clear that the limit of the Hawking mass as $t\to T^{*}(x)$, where
$T^{*}(x)\,=\,{\rm ess}\!\!\sup_{\partial M\quad}\!w_{x}^{(1)}$
@@
As already noted, a full analysis of this quantity will be deferred to future work; at present we limit ourselves to a few preliminary remarks. First, using the expansion (A.4) in Proposition A.3, we obtain
$K_{p}\,=\,\frac{\Gamma(\frac{1}{2})\,\Gamma(c_{p})}{\Gamma(a_{p}+\frac{3}{2})\Gamma(b_{p}+\frac{3}{2})}\,=\,(p-1)\,(1+o(1))\,,\qquad\hbox{as}\,\,p\to 1^{+}\,.$
If we assume an upper bound on the boundary energy of the form
@@
It follows that, within this framework, only the first integral – the so-called bulk term – contributes in the limit as $p\to 1^{+}$. Assuming the $L^{1}$-convergence of the perimeters of the sublevel sets of the $p$-IMCF to the perimeters of the sublevel sets of the IMCF with the same pole, the bulk term should converge, as $p\to 1^{+}$, to the quantity
$\mathfrak{m}_{\Lambda}^{(1)}(M,g,x)\,=\,\int_{-\infty}^{T_{\Lambda}\wedge T^{*}_{x}}\!\!\!\left(\frac{e^{\tau/2}}{16\pi}\right)\,\big(4\pi-\Lambda\,{\rm Per}(\Omega_{x,\tau}^{(1)})\big)\,d\tau\,$, (1.9)
where $T_{\Lambda}\wedge T^{*}_{x}=\min\{T_{\Lambda},T^{*}(x)\}$. This quantity, that we name Polarized $1$-harmonic Mass as it has been obtained from the Polarized $p$-harmonic Masses under a set of reasonable simplifying assumptions, is interesting in its own right and allows us to make some final considerations.
@@ 1.5. The $1$-harmonic Mass
To begin with, we note that the quantity defined in (1.9) vanishes in correspondence of the de Sitter metric
$g_{\Lambda}\,=\,\frac{dr\otimes dr}{1-\frac{\Lambda}{3}r^{2}}+r^{2}g_{\mathbb{S}^{2}}\,,\qquad 0<r\leq R_{\Lambda}=\sqrt{3/\Lambda}\,$
when computed with respect to the north pole $N$. Indeed, in this case we have that $T_{\Lambda}=T^{*}(N)$, the canonical IMCF is given by $w_{N}^{(1)}=2\log r$, and the area of its level sets is $4\pi e^{\tau}$ for every $\infty<\tau<T_{\Lambda}=2\log R_{\Lambda}$. A straightforward integration then yields
@@
It is also noteworthy that this invariant is consistent with the parameter $m$ appearing in the definition of the Schwarzschild–de Sitter metric, whose spatial part is given by
$g_{\Lambda,m}\,=\,\frac{dr\otimes dr}{1-\frac{\Lambda}{3}r^{2}-\frac{2m}{r}}+r^{2}g_{\mathbb{S}^{2}}\,,\qquad R_{\Lambda}^{-}(m)\leq r\leq R_{\Lambda}^{+}(m)\,$,
where $R_{\Lambda}^{\pm}(m)$ denote the two positive roots of the polynomial $P(r)=r-(\Lambda/3)r^{3}-2m$. By attaching a spherical cap (hemisphere) of radius $R_{\Lambda}^{-}(m)$ to the black-hole horizon of $g_{\Lambda,m}$ – that is, to the boundary component corresponding to $r=R_{\Lambda}^{-}(m)$ – so as to obtain a manifold with connected boundary and a canonical north pole, one observes that the IMCF starting from the north pole $N$ of the spherical cap and ending on the cosmological horizon $\{r=R_{\Lambda}^{+}(m)\}$ is still given by $w_{N}^{(1)}=2\log r$. The only difference lies in the global existence time of the flow, which in this case is $T^{*}(N)=T_{\Lambda}^{+}(m)=2\log R_{\Lambda}^{+}(m)$. The area of the level sets remains equal to $4\pi e^{\tau}$ for $\infty<\tau<T_{\Lambda}^{+}(m)$, and therefore the same integration as above shows that, in this setting, the value of the $1$-harmonic Mass is
$\frac{R^{+}_{\Lambda}(m)}{2}\left(1-\frac{\Lambda}{3}(R^{+}_{\Lambda})^{2}\right)=m\,.$
Having validated these consistency checks, we return to the study of compact initial data with connected boundary and note that, in this class, the $1$-harmonic mass is always nonnegative. Indeed, an IMCF with obstacle subject to our chosen normalizations exhibits subexponential growth of the areas and, more precisely, satisfies ${\rm Per}(\Omega_{x,t}^{(1)})\leq 4\pi e^{t}$, with strict inequality for $t>T_{*}(x)$ (see [73, Remark 3.5-(iii)]). A direct integration then yields
$\mathfrak{m}_{\Lambda}^{(1)}(M,g,x)\,$ $=\,\int_{-\infty}^{T_{\Lambda}\wedge T^{*}_{x}}\!\frac{e^{\tau/2}(1-\Lambda\,e^{\tau})}{4}\,d\tau\,+\Lambda\int_{-\infty}^{T_{\Lambda}\wedge T^{*}_{x}}\!\!\!\left(\frac{e^{\tau/2}}{16\pi}\right)\,\big(4\pi e^{\tau}-\,{\rm Per}(\Omega_{x,\tau}^{(1)})\big)\,d\tau\,$
$\geq\,\frac{e^{(T_{\Lambda}\wedge T^{*}_{x})/2}}{2}\,\left(1-\frac{e^{T_{\Lambda}\wedge T^{*}_{x}}}{e^{T_{\Lambda}}}\right)\,\geq\,0\,.$
Observe that if $\mathfrak{m}_{\Lambda}^{(1)}(M,g,x)=0$, then it necessarily follows from the last inequality that $T_{\Lambda}\leq T^{*}{(x)}$. Furthermore, in this situation one must also have $T_{\Lambda}\leq T_{*}(x)$; otherwise, the penultimate inequality would be strict. On the other hand, integrating equation (1.7) by parts over a domain of the form $M\setminus\Omega_{x,t}^{(1)}$ with $t<T_{*}(x)$, and employing the obstacle condition, one obtains
$4\pi\,=\,{{\rm Per}(\Omega_{x,t}^{(1)})}\,e^{-t}\,=\,\int_{\partial M}\!\!e^{-w_{x}^{(1)}}\,d\sigma\,\leq\,|\partial M|\,e^{-T_{*}(x)}\,.$
Combining the latter inequality with what was obtained previously, we find that when $\mathfrak{m}_{\Lambda}^{(1)}(M,g,x)$ vanishes, the following area bound holds
$\frac{12\pi}{\Lambda}\,=\,4\pi R_{\Lambda}^{2}\,=\,4\pi e^{T_{\Lambda}}\leq\,4\pi e^{T_{*}(x)}\leq\,|\partial M|\,.$ (1.10)
This area bound has a couple of interesting applications. The first one is a Positive Mass Theorem for initial data sets with positive Ricci curvature and an isoperimetric, totally geodesic horizon. This class of initial data was first considered by Eichmair in [33], and in fact the rigidity conclusion in the next theorem is obtained by combining (1.10) with [33, Theorem 2]:
Theorem 1.5 (Positive Mass Theorem for Isoperimetric Initial Data).
Let $(M,g)$ be a compact orientable three-dimensional Riemannian manifold with smooth boundary satisfying
${\mathrm{R}}\geq 2\Lambda>0\,\qquad\hbox{and}\qquad{\rm Ric}>0\,.$
Assume that the boundary is totally geodesic and that it is an isoperimetric surface of the doubled manifold. Then, for every $x\in M\setminus\partial M$, we have
@@
The second application of (1.10) is a classification of zero mass static metrics, inspired by the celebrated Boucher-Gibbons-Horowitz [17] inequality. Recall that a three-dimensional static triple $(M,g,V)$ is a solution to

### v2

# Mass-type invariants in the presence of a cosmological constant

@@ Abstract.
In this paper, we introduce a new family of mass-type invariants for time-symmetric initial data in space-times satisfying the Dominant Energy Condition. For positive cosmological constant, these invariants, unlike the total Hawking mass, turn out to be genuinely effective in providing new characterizations of the de Sitter solution. From a theoretical standpoint, this opens a new perspective on how one might refine the rigidity statement originally proposed by Min-Oo in his well known conjecture, later refuted by the counterexamples of Brendle, Marques, and Neves. Via a formal limiting procedure, we also define another invariant, the $1$-harmonic Mass, for which we independently prove a positive mass theorem and a Penrose-type inequality, thereby extending tools for probing space-time geometries in the presence of a positive cosmological constant.
@@
Keywords: Monotonicity formulas, $p$-Green’s function, Hawking mass, de Sitter solution, Min-Oo conjecture, Positive Mass Theorem, Penrose Inequality.
@@
The geometric intuition behind this statement is quite natural: if one postulates the validity of a positive mass theorem and assumes a strong asymptotic decay of the metric toward the model one – so that any reasonable asymptotic geometric invariant, including the yet undefined notion of mass, is forced to vanish – then it would be natural to conclude that the given manifold is isometric to the model space. In this sense, Theorem 1.2 should be regarded as a pre-Positive Mass Theorem, as it represents a preliminary condition for the validity of a full Positive Mass statement. Following a similar rationale, Min-Oo proposed in [55] his famous conjecture for initial data with a positive cosmological constant, aiming to characterize the de Sitter solution as the only one with zero mass. The precise statement of the conjecture is as follows:
@@
The general conjecture, however, does not hold under the mere natural assumptions under which it was originally formulated: Brendle, Marques and Neves [22] constructed smooth metrics on the upper hemisphere that coincide with the standard round metric in a tubular neighborhood of the equatorial boundary, having scalar curvature ${\mathrm{R}}_{g}\geq 2\Lambda>0$ everywhere and strictly grater than $2\Lambda$ on an open domain (supported far away from the round tubular neighborhood of the equator). The particular features of these counterexamples not only disprove the conjecture but also invalidate the heuristic behind it: no pre-positive mass theorem in the style of Theorem 1.2 can be obtained by imposing any order of contact with the model solution – here, the hemisphere or space-like de Sitter metric – in a neighborhood of the cosmological horizon. In physical terms, one may assert that whatever the mass associated with these initial data may be, it cannot be measured by an observer on the cosmological horizon, even under the idealized assumption that this observer possesses complete knowledge of the surroundings.
@@
We now ask whether this strategy can be implemented on a compact initial data set $(M,g)$ with connected minimal boundary and scalar curvature ${\mathrm{R}}\geq 2\Lambda>0$, in line with condition (1.2). Assume the above monotonicity formula extend rigorously to a weak setting that allows flow singularities, in the spirit of Huisken-Ilmanen [41]. Further assume there exists a weak IMCF starting from a point of the initial data set – more precisely, an ancient solution whose level sets collapse to a chosen pole as $t\to$-$\infty$ – and that this flow expires by landing on the cosmological horizon at a finite time $T^{*}$. In this thought experiment, the Hawking mass at time $T^{*}$, can be viewed as the total mass of the manifold. Monotonicity, together with $\mathfrak{m}_{\rm Haw}(\Sigma_{t})\to 0$ as $t\to$-$\infty$, then strongly suggests a Positive Mass–type statement for this total mass. This appealing line of reasoning is nevertheless ruled out by the counterexamples of Brendle, Marques, and Neves. Indeed, once the flow enters a region where ${\mathrm{R}}>2\Lambda$, the last term in (1.4) forces the derivative of the Hawking mass to be strictly positive. Thus, the Hawking mass, which vanishes at the pole and is nondecreasing along the flow, becomes strictly positive and remains so up to time $T^{*}$, when the flow expires on the cosmological horizon. In particular,
@@
However, in these counterexample-manifolds a whole neighborhood of the boundary is isometric to a tubular neighborhood of the equator in the standard hemisphere of radius $R_{\Lambda}=\sqrt{3/\Lambda}$ (equivalently, with constant sectional curvature $\Lambda/3$). A direct computation shows that the Hawking mass of this standard equator vanishes, since it is a minimal surface of area $4\pi R_{\Lambda}^{2}=12\pi/\Lambda$. Consequently, any attempt to obtain a Positive Mass-type result via such an inverse mean curvature flow faces fundamental obstructions.
Upon closer examination, it becomes clear that the underlying cause of this speculative contradiction lies in the initial assumption that a flow of the specified type actually exists. Recent progress, including the important work of Xu [74], shows that the Dirichlet problem is not the most appropriate setting in which to study the weak IMCF on a compact manifold with boundary. Instead, the natural condition appears to be that of the so‑called outer obstacle condition, in which the level sets of the arrival time function of the IMCF, namely, the weak solution of
@@
where $\nu$ denotes the outward unit normal to $\partial M$ (we refeer the reader to [74, Summary of Definitions/Theorem 1.5] for the precise definition). In general, however, the Hawking mass is not guaranteed to be nondecreasing along such a flow. A concrete counterexample is provided by the region inside a Clifford torus in the round sphere of radius $R_{\Lambda}$. This manifold has constant scalar curvature $2\Lambda>0$, so it satisfies (1.2), and its boundary is the minimal Clifford torus $\mathbb{T}_{\Lambda}$ with area $6\pi^{2}/\Lambda>12\pi/\Lambda$. A straightforward calculation then yields that the Hawking mass of this boundary is given by
@@
Now initiate the IMCF with outer obstacle $\mathbb{T}_{\Lambda}$ from a point in this manifold. As already mentioned, we have $\mathfrak{m}_{\rm Haw}(\partial\Omega_{t})\to 0$ as $t\to$-$\infty$, where $\Omega_{t}=\{w<t\}$ denotes the sublevel set. In the context of IMCF with an outer obstacle, it is more convenient to think in terms of the perimeter of the sublevels rather than the area of the level sets $\Sigma_{t}=\{w=t\}$, so as to be able to extend our considerations even beyond the first time the flow comes into contact with the boundary. As time increases, the Hawking mass of $\partial\Omega_{t}$ initially remains nonnegative by monotonicity, up to reaching a certain time $T_{0}$, smaller than the first contact time. Now, following [74], one can check that the IMCF with outer obstacle sweeps out the whole domain $M\setminus\Omega_{T_{0}}$ in finite time, say $T^{*}-T_{0}$. Moreover, for large times, the perimeter of the sublevel sets $\Omega_{t}=\{w<t\}$ approaches the full area of the boundary. In our case, this implies that for times $t$ sufficiently close to $T^{*}$, the perimeter of $\Omega_{t}$ becomes strictly greater than $12\pi/\Lambda$. As the contribution of the mean curvature term is always nonpositive, it follows that for such times the Hawking mass becomes strictly negative
$\mathfrak{m}_{\rm Haw}(\partial\Omega_{t})\leq\sqrt{\frac{{\rm Per}(\Omega_{t})}{16\pi}}\,\left\{1-\frac{\Lambda}{12\pi}{\rm Per}(\Omega_{t})\right\}\,<0.$
Thus, in this example $t\mapsto\mathfrak{m}_{\rm Haw}(\partial\Omega_{t})$ starts near 0 as $t\to$-$\infty$ and becomes stricltly negative as $t\to T^{*}$, so it cannot be nondecreasing for all times.
1.3. A potential-theoretic approach toward the Positive Mass Theorem.
@@
$K_{p}\,=\,\frac{\Gamma(\frac{1}{2})\,\Gamma(c_{p})}{\Gamma(a_{p}+\frac{3}{2})\Gamma(b_{p}+\frac{3}{2})}\,$, (1.7)
@@
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth connected boundary $\partial M$, whose scalar curvature $\mathrm{R}$} satisfies
@@
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth connected minimal boundary $\partial M$, whose scalar curvature $\mathrm{R}$} satisfies
@@
in the sense described by Xu in [74, Summary of Definitions/Theorem 1.5], $\nu(y_{0})$ being the outer unit normal at $y_{0}\in\partial M$. The normalization constant in front of the Dirac delta is reflected in the fact that, for every $t\in\mathbb{R}$,
@@
(see the proof of Theorem 3.8). Indeed, in the limit as $p\to 1^{+}$, this identity induces the following normalization for the area growth law of the IMCF
@@
Note that $T_{\Lambda}$ coincides both with the maximal existence time and with the first contact time of the model IMCF, namely the IMCF starting from the north pole $N$ of the standard sphere of radius $R_{\Lambda}$ and having the equator as an outer obstacle. Indeed, the model solution is explicitly given by $w_{N}=2\log r$ in this framework. The behavior of the structural coefficients, as $p\to 1^{+}$, makes it possible to identify a formal limit of the monotone quantities $\mathfrak{m}^{\!(p)}_{\Lambda}(x,t)$ toward the Hawking mass. In fact, assuming that the pp‑IMCF $w_{x}^{(p)}$ converge to an IMCF with obstacle $w_{x}^{(1)}$ in a way that is strong enough to justify the convergence of the areas of the level sets and of suitable integral norms of the gradients, we would obtain the following limit for small enough $t<T_{\Lambda}$:
@@
$T_{*}(x)\,=\,\min_{\partial M}w_{x}^{(1)}=\inf\,\Bigl\{t\in\mathbb{R}:\,\bigl|\partial\Omega^{(1)}_{x,t}\cap\partial M\bigr|\neq 0\Bigr\}$ (1.9)
of the weak IMCF with the obstacle $\partial M$ emanating from $x$. In this range, in fact, the standard area growth $|\Sigma_{x,t}^{(1)}|=4\pi e^{t}$ is satisfied (see [74, Remark 3.5-(iii)]) and it is very plausible that one can adapt the theory of improved convergence, recently developed by Benatti, Pluda, and Pozzetta in [14, Theorem 1.2].
On one side, the formal convergence to the Hawking mass for fixed times clarifies the geometric nature of our monotone quantities. On the other side, it is already clear that the limit of the Hawking mass as $t\to T^{*}(x)$, where
$T^{*}(x)\,=\,\max_{\partial M}\,w_{x}^{(1)}$ (1.10)
@@
As already noted, a full analysis of this quantity will be deferred to future work; at present we limit ourselves to a few preliminary remarks. First, using expansion (A.4) in Proposition A.3, we obtain
$K_{p}\,=\,(p-1)\,(1+o(1))\,,\qquad\hbox{as}\,\,p\to 1^{+}\,$,
where $K_{p}$ is the constant defined in (1.7). If we assume an upper bound on the boundary energy of the form
@@
It follows that, within this framework, only the first integral – the so-called bulk term – contributes in the limit as $p\to 1^{+}$. Assuming the $L^{1}$-convergence of the perimeters of the sublevel sets of the $p$-IMCF to the perimeters of the sublevel sets of the IMCF with obstacle with the same pole, the bulk term should converge, as $p\to 1^{+}$, to the quantity
$\mathfrak{m}_{\Lambda}^{(1)}(M,g,x)\,=\,\int_{-\infty}^{T^{+}_{x}}\!\!\!\left(\frac{e^{\tau/2}}{16\pi}\right)\,\big(4\pi-\Lambda\,{\rm Per}(\Omega_{x,\tau}^{(1)})\big)\,d\tau\,$, (1.11)
where $T^{+}_{x}=\min\{T_{\Lambda},T^{*}(x)\}$. This quantity, that we name Polarized $1$-harmonic Mass as it has been obtained in the limit as $p\to 1^{+}$ of the Polarized $p$-harmonic Masses under a set of reasonable simplifying assumptions, is interesting in its own right and allows us to introduce a second group of main results related to it. Notably, these include a Positive Mass Theorem and a Riemannian Penrose Inequality for the $1$-harmonic mass.
@@ 1.5. The $1$-harmonic Mass
We first observe that, on a compact manifold with smooth boundary, the quantity (1.11) is well defined for every $x\in M\setminus\partial M$. Its definition indeed only requires the existence of an IMCF with outer obstacle $\partial M$ emanating from the pole $x$, subject to a suitable normalization, that is, a solution of (7.1). Once such a flow is available, one can indeed compute the areas of its sublevel sets and the last contact time (1.10) of the flow with the obstacle, denoted by $T^{*}(x)$. This time, when compared with the standard existence time $T_{\Lambda}$, uniquely determines the upper limit of integration $T^{+}_{x}$. We also note that the polarized $1$-harmonic mass (1.11) vanishes in the correspondence of the de Sitter metric
$g_{\Lambda}\,=\,\frac{dr\otimes dr}{1-\frac{\Lambda}{3}r^{2}}+r^{2}g_{\mathbb{S}^{2}}\,,\qquad 0<r\leq R_{\Lambda}=\sqrt{3/\Lambda}\,$,
when computed with respect to the north pole $N$. Indeed, in this case we have that $T_{\Lambda}=T^{*}(N)$, the canonical IMCF is given by $w_{N}^{(1)}=2\log r$, and the area of its level sets is $4\pi e^{\tau}$ for every $\infty<\tau\leq T_{\Lambda}=2\log R_{\Lambda}$. A direct evaluation of the defining integral then yields
@@
In the next theorem, we prove that, for a suitable class of time-symmetric initial data, the polarized $1$-harmonic Mass is always nonnegative, and that its vanishing characterizes the de Sitter solution.
Theorem 1.5 (Positive Mass Theorem for the Polarized $1$-harmonic Mass).
Let $(M,g)$ be a compact three-dimensional Riemannian manifold with smooth minimal boundary $\partial M$, whose scalar curvature satisfies
${\mathrm{R}}\geq 2\Lambda,$
for some $\Lambda>0$. Assume that $\partial M$ is simply connected and unstable and that there are no closed minimal surfaces in $M\setminus\partial M$. Then, for every $x\in M\setminus\partial M$, we have
@@
Remark 1.6.
As it will be clear from Lemma 7.1 below, the absence of closed minimal surfaces in $M\setminus\partial M$, in the presence of a simply connected and unstable minimal boundary, implies that the manifold is topologically simple and in fact diffeomorphic to $\mathbb{B}^{3}$. This property is important for the rigidity statement, but it is irrelevant for the mere inequality, whose proof does not rely on any topological assumption.
Remark 1.7.
We conjecture that following the ideas developed in Section 6 a similar positive mass statement can be proven for the unpolarized version of the $1$-harmonic Mass, namely
$\mathfrak{m}_{\Lambda}^{(1)}(M,g)\,=\,\inf_{x\in M}\,\mathfrak{m}_{\Lambda}^{(1)}(M,g,x)\,.$
In fact, a closer inspection of formula (1.11) suggests that when the pole approaches the boundary the value of the Polarized $1$-harmonic Mass would increase.
To introduce our last main result, we first note that the $1$-harmonic mass defined in (1.15) is compatible with the real parameter $m$ appearing in the two-parameter family of Schwarzschild-de Sitter metrics, whose spatial component we now describe. These Riemannian metrics are given by
$g_{\Lambda,m}\,=\,\frac{dr\otimes dr}{1-\frac{\Lambda}{3}r^{2}-\frac{2m}{r}}+r^{2}g_{\mathbb{S}^{2}}\,,\qquad R_{\Lambda}^{-}(m)\leq r\leq R_{\Lambda}^{+}(m)\,$, (1.12)
with $\Lambda>0$ and $0\leq m<1/(3\sqrt{\Lambda})$. Here $R_{\Lambda}^{\pm}(m)$ denote the two positive roots of the equation $p_{\Lambda}(r)=m$, where $r\mapsto p_{\Lambda}(r)$ is the cubic polynomial
$p_{\Lambda}(r)\,=\,\frac{r}{2}\left(1-\frac{\Lambda}{3}r^{2}\right).$
For future convenience, we also define $T_{\Lambda}^{\pm}(m)=2\log(R_{\Lambda}^{\pm}(m))$ and observe that these two numbers represent the endpoints of the standard existence interval of the model IMCF $w=2\log r$. In other words, this flow originates from the black hole horizon $\{r=R_{\Lambda}^{-}(m)\}$ at time $T_{\Lambda}^{-}(m)$ and reaches the cosmological horizon $\{r=R_{\Lambda}^{+}(m)\}$ at time $T_{\Lambda}^{+}(m)$. Furthermore, note that for $\Lambda>0$ the polynomial $r\mapsto p_{\Lambda}(r)$ is increasing on the interval $R_{\Lambda}^{-}(0)=0\leq r\leq\sqrt{1/\Lambda}$ and decreasing on the interval $\sqrt{1/\Lambda}\leq r\leq\sqrt{3/\Lambda}=R_{\Lambda}^{+}(0)$. In particular, for every $\Lambda>0$ and every $0\leq m<1/(3\sqrt{\Lambda})$, we have
$0\leq R_{\Lambda}^{-}(m)<\sqrt{{1}/{\Lambda}}<R_{\Lambda}^{+}(m)\leq R_{\Lambda}\,.$